estimate the radius r, of a circular region with an area of 27ft^2 where A=pi*r^2
r=? ft
I tried to use 27ft^2=729ft pi=3.14 , but don't know how to solve for r in feet? Could you please explain to me what I should do? Thanks!

Answer by Alan3354(21583)

Area = pi*r^2
27 = pi*r^2
r^2 = 27/pi = ~8.5944
r =~ 2.932 feet